# May the Feedback Be with You! Unlocking the Power of Feedback-Driven Deep Learning Framework Fuzzing via LLMs

### v3

@@ Abstract.
To fill this gap, we propose FUEL to effectively utilize the feedback information, which comprises two Large Language Models (LLMs): analysis LLM and generation LLM. Specifically, analysis LLM infers analysis summaries from feedback information, while the generation LLM creates tests guided by these summaries. Furthermore, based on multiple feedback guidance, we design two additional components: (i) a feedback-aware simulated annealing algorithm to select operators for test generation, enriching test diversity. (ii) a program self-repair strategy to automatically repair invalid tests, enhancing test validity. We evaluate FUEL on the two most popular DL frameworks, and experiment results show that FUEL can improve line code coverage of PyTorch and TensorFlow by 9.15% and 14.70% over state-of-the-art baselines (e.g., TitanFuzz and WhiteFox). By the time of submission, FUEL has detected 104 previously unknown bugs for PyTorch and TensorFlow, with 93 confirmed as new bugs, 49 already fixed, and 14 assigned CVE IDs.
copyright: nonejournalyear: 2025
1. Introduction
@@
Deep Learning (DL) techniques have achieved success in various domains over the last decade, including healthcare (Brown, 2021), chatbots (OpenAI, 2023a), autonomous driving (Huval et al., 2015), and finance security (Tugend, 2021). In these domains, popular DL frameworks like PyTorch (PyTorch, 2018) and TensorFlow (TensorFlow, 2020) help developers build, train, compile, and deploy DL models. Despite their widespread use, these frameworks remain susceptible to bugs, which can pose serious consequences, particularly in safety-critical applications (ube, 2021). Hence, ensuring the quality of DL frameworks is essential, and various automated testing techniques (Wang et al., 2020; Pham et al., 2019; Mu et al., 2024; Liu et al., 2023a; Dutta et al., 2021) are proposed to detect bugs in DL frameworks. Among them, fuzzing (Sutton et al., 2007; Herrera et al., 2021; Gong et al., 2025) is a prominent technique that systematically generates numerous inputs to uncover potential bugs. Traditional fuzzers (Wang et al., 2020; Liu et al., 2023a) typically require manual design for mutation rules or operator specifications, which demands extensive domain knowledge and significant human effort.
@@
To alleviate the above issue, recent research has introduced Large Language Models (LLMs) (OpenAI, 2023a, b) into fuzzing, motivated by their success in other fields (Fan et al., 2023; Li et al., 2024). One of the first classical studies is TitanFuzz (Deng et al., 2023), whose key insight is that LLMs can implicitly learn both language syntax and intricate DL API constraints for DL program generation. Inspired by TitanFuzz, some follow-up studies (Deng et al., 2024b; Lyu et al., 2024; Yang et al., 2024, 2025) are proposed to explore the potential of LLMs in fuzzing. The core idea behind these studies is to leverage the strong generative abilities of LLMs to generate or mutate inputs directly. Under such a trend, the “magnificent edifice” of fuzzing has already been built, but one dark cloud still hangs in the air: Have we fully exploited the potential of feedback information during the fuzzing loop, especially the semantically rich ones? Before answering this question, we first analyze the limitations of existing studies:
Figure 1. Motivation example of illustrating Limitation #1&2
@@
Limitation #1: Overlooking multiple types of feedback. Previous approaches primarily focus on leveraging the powerful generation capabilities of LLMs to generate random inputs or edge cases. However, they tend to overlook the rich information derived from multiple feedback types. For example, TitanFuzz (Deng et al., 2023) and FuzzGPT (Deng et al., 2024b) do not utilize any feedback in their fuzzing loop, merely employing differential testing (Evans and Savoia, 2007) to identify potential bugs. Regrettably, the absence of feedback guidance leads to decreased code coverage and leaves more bugs undetected. In contrast, PromptFuzz (Lyu et al., 2024) employs a coverage-guided strategy to encourage LLMs to generate more diverse tests, alleviating the coverage plateaus problem. WhiteFox (Yang et al., 2024) advances this idea by injecting compiler-optimization heuristics into the feedback loop. However, these techniques still only utilize a single and simple type of feedback during the whole fuzzing campaign and fail to consider multiple types of feedback simultaneously. As Figure 1 shows, each test execution yields three mutually exclusive feedback categories: coverage information, bug reports, and exception logs. However, how to fuse and exploit such complex feedback remains unexplored.
@@
Limitation #2: Coarse-grained feedback information analysis. Each test execution in fuzzing yields various semantic information, which is usually a mixture of programming language and natural language. state-of-the-art (SOTA) feedback-driven fuzzers still distill feedback information into coarse signals (e.g., coverage increase), thereby discarding the fine-grained, context-rich semantic information. Even PromptFuzz (Lyu et al., 2024), which mitigates coverage plateaus, only focuses on whether new code is covered, ignoring semantics from concrete code snippets. This coarse-grained manner prevents the fuzzer from reasoning about the actual runtime state of the system under test (SUT). As Figure 1 shows, coverage information pinpoints specific code snippets, bug reports contain symptom descriptions, and exception logs provide detailed stack traces. Each feedback category can be effectively processed by LLMs, which transform raw feedback into actionable insights to guide subsequent test generation.
@@
Figure 2. Motivation example of illustrating Limitation #3
Limitation #3: Insufficient test diversity. TitanFuzz and FuzzGPT restrict themselves to API-level fuzzing (introduced in § 2.1), exhaustively enumerating isolated operators one by one. Although this paradigm achieves high API coverage, it overlooks bugs that manifest only when multiple operators interact within DL models. Model-level fuzzers such as WhiteFox (Yang et al., 2024) synthesize tests that target compiler optimization patterns (e.g., operator fusion), yet this optimization-centric focus sacrifices test diversity, as many latent bugs lie outside the chosen optimization patterns. For instance, Figure 2 shows a test in which torch.frexp returns two tensors (torch.int32 and torch.float32) which are subsequently element-wise multiplied. Eager backend of PyTorch silently handles type promotion, whereas the compiler backend crashes with CppCompileError. TitanFuzz misses this bug due to their single-API testing paradigm. WhiteFox was also unable to detect this bug, as it can not generate such models that mismatch any optimization pattern. A natural and intuitive idea is to make LLM consider more combinations of operators through retrieval-augmented generation (RAG) (Gao et al., 2023), diversifying generated tests.
Limitation #4: Limited validity of generated tests.
Figure 3. Motivation example of illustrating Limitation #4
Even SOTA LLMs currently struggle to create valid tests for DL frameworks (Deng et al., 2024a). Existing techniques prompt LLMs with few-shot examples through in-context learning, but this paradigm poses limitations to LLMs’ understanding of the constraints on DL operators. Figure 3 shows an invalid test generated by WhiteFox that combines torch.mm and Add, imposing two constraints: (i) the first input of torch.mm must be column-aligned with the second; (ii) the inputs of Add must be either identically shaped or broadcastable. The generated tensors, however, have shapes $(2\times 10)$ and $(2\times 10)$, violating above constraints and triggering a shape-mismatch error, which is a false positive. Our manual check reveals that 47.54% (290/610) of the tests generated by WhiteFox are invalid, similar to the above example. These failures not only waste fuzzing budgets but also lose opportunities to uncover bugs within operator interactions. Notably, we observe that most invalid tests are easy to fix. In this case, simply reshaping x1 from $(2\times 10)$ to $(2\times 2)$ generates a valid test that executes successfully and explores previously uncovered code paths. Equipped with the exception logs, LLMs are capable of automatically conducting such small-scale repairs, converting invalid executions into valid ones, thereby improving code coverage and bug detection capabilities.
@@
Our work. In this paper, we propose FUEL, the first feedback-driven fuzzer that fully utilizes feedback information with two LLM-based agents, namely analysis LLM ($\mathcal{M_{A}}$) and generation LLM ($\mathcal{M_{G}}$). The key idea of our design is that LLMs could understand and analyze valuable feedback information. $\mathcal{M_{A}}$ first automatically infers the analysis summaries from the feedback. Then, analysis summaries are used to guide the $\mathcal{M_{G}}$ to generate tests that help to explore new code coverage or trigger new bugs (addressing Limitation #1 and Limitation #2). However, tests generated by LLMs usually contain common well-tested operators (e.g., Conv2d, ReLU), resulting in the lack of some uncommon poorly-tested operators (e.g., ConvTranspose, InstanceNorm). This behavior reduces test diversity. To solve this, we collect core operators from the DL framework API document as a retrieval set and design a feedback-aware simulated annealing algorithm to select operators from the retrieval set in each iteration. The retrieved operators, along with the analysis summary, are filled into the prompt of the $\mathcal{M_{G}}$ as context, allowing $\mathcal{M_{G}}$ to generate tests by using more operators (addressing Limitation #3). To improve test validity, FUEL introduces a program self-repair strategy to repair invalid tests. In simple terms, $\mathcal{M_{A}}$ in FUEL generates the repair strategy based on the exception log, which prompts the $\mathcal{M_{G}}$ to repair the test from the last iteration (addressing Limitation #4).
@@
  • Evaluation. Our experimental results show that the tests generated by FUEL achieve superior performance compared to SOTA baselines (TitanFuzz and WhiteFox), improving line coverage on two DL frameworks (PyTorch and TensorFlow) by 9.15% and 14.70%, respectively. Additionally, ablation studies validate the contribution and effectiveness of each component in FUEL ’s design.
@@
  • Real-World Contribution. So far, FUEL has detected 104 previously unknown bugs, with 93 already confirmed, 49 already fixed, and 14 assigned CVE IDs. 14 detected bugs were labeled as high priority, and one issue was labeled as utmost priority.
2. Background
Figure 4. Overview of FUEL
2.1. DL Framework Testing
@@
Previous studies on DL framework testing can be classified into API-level fuzzers (Wei et al., 2022a; Xie et al., 2022; Deng et al., 2023) and model-level fuzzers (Wang et al., 2020; Liu et al., 2023a; Yang et al., 2024). API-level fuzzers target testing each DL framework API by infilling its parameters. Notably, a DL framework API usually corresponds to a DL operator. For instance, torch.Conv2D, a PyTorch API (corresponding to 2-D convolution operator) can be filled with some parameters by API-level fuzzers, including kernel_size, stride, and padding. Then, a 4-D tensor will be created as input for torch.Conv2D. Lastly, comparing the outputs on different hardware (e.g., CPU and GPU) to detect the potential bugs.
@@
However, this paradigm is limited in its ability to detect bugs resulting from the intricate compositions of diverse operators in DL models. Therefore, model-level fuzzers are proposed to synthesize or mutate DL models that encapsulate a wide range of operators. A model-level test typically comprises two components: model definition and input tensors. As shown in Figure 3, the model definition is a Python class inheriting from the framework’s base Model class (e.g., torch.nn.Module). Meanwhile, the input tensors supply external tensor data to the model (e.g., torch.randn(4, 1)). Model-level fuzzers therefore strive to generate both DL model (i.e., model definition) and input tensors that satisfy the model constraint, jointly forming a single model-level test. In this work, our designed FUEL is a model-level fuzzer, which aims to create diverse yet valid DL models.
2.2. LLM-based Fuzzing
@@
Until now, LLMs have achieved remarkable success in code intelligence tasks (Hou et al., 2024). Modern LLMs, typically based on Transformers, can be categorized into two types: Dense Transformer (Vaswani et al., 2017) and Mixture-of-Experts (MoE) architecture (Jiang et al., 2024). Dense Transformers (e.g., Llama-3 (Dubey et al., 2024) and Qwen-2.5-Coder (Hui et al., 2024)) employ uniform parameter activation across all layers, enabling full contextual interaction through self-attention mechanisms, while MoE architecture models (e.g., Mixtral (Jiang et al., 2024) and DeepSeek-V3 (DeepSeek-AI, 2024)) introduce sparse activation via expert routing algorithms, dynamically allocating computations to specialized subnetworks for improved scalability and efficiency. Recently, reasoning models (e.g., DeepSeek-R1 (DeepSeek-AI, 2025), GPT-o1 (ope, 2024)) based on chain-of-thought (CoT) (Wei et al., 2022b) have attracted worldwide attention, showing exciting performance on math- and code-related benchmarks.
@@
Impressive performance of LLMs motivates researchers to bring LLMs into software testing (Wang et al., 2024a), especially fuzzing (Huang et al., 2024). In LLM-based fuzzing, LLM is usually employed as the test generator. Along with instructions and few-shot examples in prompts, it directly generates test code (similar to the code snippet in Figure 3). Unfortunately, relying on a single generation LLM within the fuzzing loop limits both the test diversity and validity. In this context, FUEL augments the fuzzing loop with an additional analysis LLM that extracts and analyzes actionable feedback information, thereby enriching test diversity and validity.

### v4

# May the Feedback Be with You! Breaking the Seal of Feedback-Driven Deep Learning Framework Fuzzing via Large Language Models

@@ Abstract
To fill this gap, we propose FUEL to effectively utilize the feedback information, which comprises two Large Language Models (LLMs): analysis LLM and generation LLM. Specifically, analysis LLM infers analysis summaries from feedback information, while the generation LLM creates tests guided by these summaries. Furthermore, based on multiple feedback guidance, we design two additional components: (i) a feedback-aware simulated annealing algorithm to select operators for test generation, enriching test diversity. (ii) a program self-repair strategy to automatically repair invalid tests, enhancing test validity. We evaluate FUEL on the two most popular DL frameworks, and experiment results show that FUEL can improve line code coverage of PyTorch and TensorFlow by 4.48% and 9.14% over four state-of-the-art baselines. By the time of submission, FUEL has detected 104 previously unknown bugs for PyTorch and TensorFlow, with 93 confirmed as new bugs, 53 already fixed. 14 vulnerabilities have been assigned CVE IDs, among which 7 are rated as high-severity with a CVSS score of "7.5 HIGH".
1 Introduction
@@
Deep Learning (DL) techniques have achieved success in various domains over the last decade, including healthcare [healthcare], chatbots [chatgpt], autonomous driving [huval2015autonomousdriving], and finance security [finance]. In these domains, popular DL frameworks like PyTorch [PyTorch] and TensorFlow [Tensorflow] help developers build, train, compile, and deploy DL models. Despite their widespread use, these frameworks remain susceptible to bugs, which can pose serious consequences, particularly in safety-critical applications [uberkill]. Hence, ensuring the quality of DL frameworks is essential, and various automated testing techniques [wang2020lemon, pham2019cradle, mu2024devmut, liu2023nnsmith] are proposed to detect bugs in DL frameworks. Among them, fuzzing [SuttonFuzzingBook, herrera2021seed, gong2025snowplow] is a prominent technique that systematically generates numerous inputs to uncover potential bugs. Traditional fuzzers [wang2020lemon, liu2023nnsmith] typically require manual design for mutation rules or operator specifications, which demands extensive domain knowledge and great human effort.
@@
To alleviate the above issue, recent research has introduced Large Language Models (LLMs) [chatgpt, openai2023gpt4] into fuzzing, motivated by their success in other fields [SEsur]. One of the first classical studies is TitanFuzz [titanfuzz], whose key insight is that LLMs can implicitly learn both language syntax and intricate DL API constraints for DL program generation. Inspired by TitanFuzz, some follow-up studies [fuzzgpt, promptfuzz, whitefox, kernelgpt] are proposed to explore the potential of LLMs in fuzzing. The core idea behind these studies is to leverage the strong generative abilities of LLMs to generate or mutate inputs directly. Under such a trend, the “magnificent edifice” of fuzzing has already been built, but one dark cloud still hangs in the air: Have we fully exploited the potential of feedback information during the fuzzing loop, especially the semantically rich ones? Before answering this question, we first analyze the limitations of existing studies:
Figure 1: Motivation example of illustrating Limitation #1&2
@@
Limitation #1: Overlooking multiple types of feedback. Previous approaches primarily focus on leveraging the powerful generation capabilities of LLMs to generate random inputs or edge cases. However, they tend to overlook the rich information derived from multiple feedback types. For example, TitanFuzz [titanfuzz] and FuzzGPT [fuzzgpt] do not utilize any feedback in their fuzzing loop, merely employing differential testing [evans2007difftesting] to identify potential bugs. Regrettably, the absence of feedback guidance leads to decreased code coverage and leaves more bugs undetected. In contrast, PromptFuzz [promptfuzz] employs a coverage-guided strategy to encourage LLMs to generate more diverse tests, alleviating the coverage plateaus problem. WhiteFox [whitefox] advances this idea by injecting compiler-optimization heuristics into the feedback loop. However, these techniques still only utilize a single and simple type of feedback during the whole fuzzing campaign and fail to consider multiple types of feedback simultaneously. As Figure 1 shows, each test execution yields three mutually exclusive feedback categories: coverage information, bug reports, and exception logs. However, how to fuse and exploit such complex feedback remains unexplored.
@@
Limitation #2: Coarse-grained feedback information analysis. Each test execution in fuzzing yields various semantic information, which is usually a mixture of programming language and natural language. state-of-the-art (SOTA) feedback-driven fuzzers still distill feedback information into coarse signals (e.g., coverage increase), thereby discarding the fine-grained, context-rich semantic information. Even PromptFuzz [promptfuzz], which mitigates coverage plateaus, only focuses on whether new code is covered, ignoring semantics from concrete code snippets. This coarse-grained manner prevents the fuzzer from reasoning about the actual runtime state of the system under test (SUT). As Figure 1 shows, coverage information pinpoints specific code snippets, bug reports contain symptom descriptions, and exception logs provide detailed stack traces. Each feedback category can be effectively processed by LLMs, which transform raw feedback into actionable insights to guide subsequent test generation.
@@
Figure 2: Motivation example of illustrating Limitation #3
Limitation #3: Insufficient test diversity. TitanFuzz and FuzzGPT restrict themselves to API-level fuzzing (introduced in § 2.1), exhaustively enumerating isolated operators one by one. Although this paradigm achieves high API coverage, it overlooks bugs that manifest only when multiple operators interact within DL models. Model-level fuzzers such as WhiteFox [whitefox] synthesize tests that target compiler optimization patterns (e.g., operator fusion), yet this optimization-centric focus sacrifices test diversity, as many latent bugs lie outside the chosen optimization patterns. For instance, Figure 2 shows a test in which torch.frexp returns two tensors (torch.int32 and torch.float32) which are subsequently element-wise multiplied. Eager backend of PyTorch silently handles type promotion, whereas the compiler backend crashes with CppCompileError. TitanFuzz misses this bug due to their single-API testing paradigm. WhiteFox was also unable to detect this bug, as it can not generate such models that mismatch any optimization pattern. A natural and intuitive solution is to enable the LLM to consider more operator combinations by supplying external context information, thereby diversifying the generated tests.
Figure 3: Motivation example of illustrating Limitation #4
Limitation #4: Limited validity of generated tests. Even SOTA LLMs currently struggle to create valid tests for DL frameworks [deng2024can]. Existing techniques prompt LLMs with few-shot examples through in-context learning, but this paradigm poses limitations to LLMs’ understanding of the constraints on DL operators. Figure 3 shows an invalid test generated by WhiteFox that combines torch.mm and Add, imposing two constraints: (i) the first input of torch.mm must be column-aligned with the second; (ii) the inputs of Add must be either identically shaped or broadcastable. The generated tensors, however, have shapes $(2\times 10)$ and $(2\times 10)$, violating the above constraints and triggering a shape-mismatch error, which is a false positive. Our manual check reveals that 47.54% (290/610) of the tests generated by WhiteFox are invalid, similar to the above example. These failures not only waste fuzzing budgets but also lose opportunities to uncover bugs within operator interactions. Notably, we observe that most invalid tests are easy to fix. In this case, simply reshaping x1 from $(2\times 10)$ to $(2\times 2)$ generates a valid test that executes successfully and explores previously uncovered code paths. Equipped with the exception logs, LLMs are capable of automatically conducting such small-scale repairs, converting invalid executions into valid ones, thereby improving code coverage and bug detection capabilities.
@@
Our work. In this study, we propose FUEL, the first feedback-driven fuzzer that fully utilizes feedback information with two LLM-based agents, namely analysis LLM ($\mathcal{M_{A}}$) and generation LLM ($\mathcal{M_{G}}$). The key idea of our design is that LLMs could understand and analyze valuable feedback information. $\mathcal{M_{A}}$ first automatically infers the analysis summaries from the feedback. Then, analysis summaries are used to guide the $\mathcal{M_{G}}$ to generate tests that help to explore new code coverage or trigger new bugs (addressing Limitation #1 and Limitation #2). However, tests generated by LLMs usually contain common well-tested operators (e.g., Conv2d, ReLU), resulting in the lack of some uncommon poorly-tested operators (e.g., ConvTranspose, InstanceNorm). This behavior reduces test diversity. To solve this, we collect core operators from the DL framework API document as a retrieval set and design a feedback-aware simulated annealing algorithm to select operators from the retrieval set in each iteration. The retrieved operators, along with the analysis summary, are filled into the prompt of the $\mathcal{M_{G}}$ as context, allowing $\mathcal{M_{G}}$ to generate tests by using more operators (addressing Limitation #3). To improve test validity, FUEL introduces a program self-repair strategy to repair invalid tests. In simple terms, $\mathcal{M_{A}}$ in FUEL generates the repair strategy based on the exception log, which prompts the $\mathcal{M_{G}}$ to repair the test from the last iteration (addressing Limitation #4).
@@
  • Evaluation. Our experimental results show that the tests generated by FUEL achieve superior performance compared to SOTA baselines (NeuRI, TitanFuzz, NNSmith, and WhiteFox), improving line coverage on two DL frameworks (PyTorch and TensorFlow) by 4.48% and 9.14%, respectively. Additionally, ablation studies validate the contribution and effectiveness of each component in FUEL ’s design.
@@
  • Real-World Contribution. So far, FUEL has detected 104 previously unknown bugs, with 93 already confirmed, 53 already fixed, and 14 assigned CVE IDs. 14 detected bugs were labeled as high priority, and one issue was labeled as utmost priority.
2 Background
2.1 DL Framework Testing
@@
Previous studies on DL framework testing can be classified into API-level fuzzers [wei2022free, xie2022docter, titanfuzz] and model-level fuzzers [wang2020lemon, liu2023nnsmith, whitefox]. API-level fuzzers target testing each DL framework API by infilling its parameters. Notably, a DL framework API usually corresponds to a DL operator. For instance, torch.Conv2D, a PyTorch API (corresponding to 2-D convolution operator) can be filled with some parameters by API-level fuzzers, including kernel_size, stride, and padding. Then, a 4-D tensor will be created as input for torch.Conv2D. Finally, comparing outputs across different hardware (e.g., CPU and GPU) to detect potential bugs.
@@
However, this paradigm is limited in its ability to detect bugs resulting from the intricate compositions of diverse operators in DL models. Therefore, model-level fuzzers are proposed to synthesize or mutate DL models that encapsulate a wide range of operators. A model-level test typically comprises two components: model definition and input tensors. As shown in Figure 3, the model definition is a Python class inheriting from the framework’s base Model class (e.g., torch.nn.Module). Meanwhile, the input tensors supply external tensor data to the model (e.g., torch.randn(4, 1)). Model-level fuzzers therefore strive to generate both the DL model (i.e., model definition) and input tensors that satisfy the model constraint, jointly forming a single model-level test. In this work, our designed FUEL is a model-level fuzzer, which aims to create diverse yet valid DL models.
Figure 4: Overview of FUEL
2.2 LLM-based Fuzzing
@@
Until now, LLMs have achieved remarkable success in code intelligence tasks [hou2024llm4se]. Modern LLMs, typically based on Transformers, can be categorized into two types: Dense Transformer [transformer] and Mixture-of-Experts (MoE) architecture [jiang2024mixtral]. Dense Transformers (e.g., Llama-3 [dubey2024llama] and Qwen-2.5-Coder [hui2024qwen2.5-coder]) employ uniform parameter activation across all layers, enabling full contextual interaction through self-attention mechanisms, while MoE architecture models (e.g., Mixtral [jiang2024mixtral] and DeepSeek-V3 [deepseekai2024deepseekv3]) introduce sparse activation via expert routing algorithms, dynamically allocating computations to specialized subnetworks for improved scalability and efficiency. Recently, reasoning models (e.g., DeepSeek-R1 [deepseekai2025deepseekr1], GPT-o1 [openai-o1]) based on chain-of-thought (CoT) [wei2022cot] have attracted worldwide attention, showing impressive performance on math- and code-related benchmarks.
@@
Impressive performance of LLMs motivates researchers to bring LLMs into software testing [wang2024software], especially fuzzing [llmfuzzsurvey]. In LLM-based fuzzing, LLM is usually employed as the test generator to synthesize programs. Along with instructions and the few-shot examples in prompts, it directly generates test code (similar to the code snippet in Figure 3). Specifically, these approaches primarily focus on using LLMs to generate or mutate inputs based on static prompts. However, they often overlook the execution results of generated tests, failing to utilize such valuable feedback to guide subsequent generation. In this context, FUEL augments the fuzzing loop with an analysis LLM to distill raw feedback information into actionable summaries, explicitly guiding the generation LLM to enrich test diversity and validity.
3 FUEL Design
Figure 4 depicts the end-to-end workflow of FUEL, integrating a $\mathcal{M_{A}}$ with a feedback-aware simulated annealing algorithm to utilize feedback information in a fully automated manner and subsequently combining analysis results to guide $\mathcal{M_{G}}$ to generate diverse and valid DL models. We show the details of each phase in the following subsections.
3.1 Prompting LLM for Test Generation
To accommodate these two scenarios, we design specialized default generation prompts that ensure continuous test generation even in scenarios where feedback is unavailable or repair stagnation occurs. The transition between feedback-driven and default generation is algorithmically controlled to optimize exploration-exploitation balance throughout the whole fuzzing campaign. Figure 5(a) shows a few-shot example in the default generation prompt. Firstly, it lists some operators selected by the heuristic search algorithm (discussed in § 3.4). Then this example goes with the instruction “Please generate a valid PyTorch model with selected operators”. Next, it provides a DL model consisting of two components: model definition and input tensors (introduced in § 2.1). Finally, with such few-shot examples, $\mathcal{M_{G}}$ generates tests with the prompt template shown in Figure 5(b). Placeholder [LIBRARY] in the instruction denotes the specific framework (PyTorch or TensorFlow), [SELECTED_OPS] is filled by results of heuristic search, and [PLACEHOLDER...] is the DL model to be generated by $\mathcal{M_{G}}$.